OpenStudy (anonymous):

Reduce ech ration to it's simplest form. 14:21 and 8:16

OpenStudy (anonymous):

each*

OpenStudy (anonymous):

each*

OpenStudy (anonymous):

Do you know primefactorization?

OpenStudy (anonymous):

Yeah, I do.

OpenStudy (anonymous):

Then primefactor both sides and divide by the common factors

OpenStudy (anonymous):

\[14= 2*7\] \[21=3*7\] What's the largest number both sides can be divided by?